### v1

@@ 5. Oblivious Authentication
As a foundation for Seldom, we design an oblivious authentication protocol for onion routing. It authenticates a client to an exit relay without revealing its identity. Clients have to obtain temporal IP certificates (Section 5.1) before being able to establish a circuit (Section 5.2) such that exit relays are guaranteed that any outgoing traffic can be irrefutably linked to the client (Section 5.3). In Appendix A, we detail how we integrate this protocol with the existing Tor handshake.
@@ 5.1. Temporal IP Certificates
Seldom needs a way to reliably represent an individual’s identity if deanonymization is required later on. Regular, CA-issued certificates are not applicable to Seldom as (i) obtaining them is costly and not always possible (Backes et al., 2014) and (ii) they would share excessive personal information with entry relays. Instead, we propose to identify users via temporal IP certificates, which link a client’s ephemeral public key to a timestamped IP address. While not necessarily allowing to reveal a criminal’s true identity directly, temporal IP certificates mimic the information LEA would obtain if the criminal could not make use of an anonymity network, thereby supporting further investigation. Only the entry relay obtains the temporal IP certificate in the clear for validation purposes. If the deanonymization reveals an invalid certificate, this serves as an indication for LEA that the client and entry relay colluded, pointing to the entry relay for further investigation.
Temporal IP certificates are issued in an automated manner by one or more CAs that are trusted by LEA. The client generates a public-private key pair for any new Internet connection (e.g., ISP-triggered reassignment of IPv4 addresses) and sends the public key to a CA for certification. In the simplest case, the CA extracts the client’s IP address from the request and creates the temporal IP certificate from that IP address, the public key, and the current timestamp. We further discuss adaptions to this process for users relying on bridge relays, e.g., for censorship evasion, in Appendix B.
@@ 5.2.1. Authentication to the Entry Relay
The purpose of our authentication protocol is to bind a temporal IP certificate to the hops within a circuit such that relays can only obtain information related to their direct predecessor. We hence refer to the keying material required to protect individual hops as hop keys from now on.
After obtaining a valid temporal IP certificate (cf. Section 5.1), the client confidentially sends that certificate to the entry relay , which extracts its public key as the first hop key . The entry relay validates by verifying the certificate’s signature by the CA and ensuring that the client uses the certified IP address. Failures to validate the IP address will surface during deanonymizations; hence, is incentivized to perform this check honestly. Next, the client generates a second hop key to prepare the authentication forwarding between and . By signing using the validated , can in turn validate the authenticity of . Finally, threshold-encrypts the client’s IP certificate, , and the signature under such that only the consortium can access that data from now on, yielding the client’s encrypted identity from the perspective of . Our hybrid threshold-encrpytion scheme will be open-sourced and is further explained in Appendix C.
@@ 5.2.2. Oblivious Authentication Forwarding
Assuming that an obliviously authenticated partial circuit has been established up to relay , we now describe how to extend the authentication via another hop to . On a high level, forwards its view on the state of current authentication, and as well as its own identity to . is a public key and usually published via the relay’s descriptor (Dingledine et al., 2004). Other relays, such as unlisted bridge relays, can be identified pseudonymously via additional, Seldom-specific certificates. Only in cases where LEA must identify a relay operator, bridge authorities would reveal the certificate owner. further signs the transmitted information under its own identity, yielding , which is verified by . Verifying the identity of reassures that LEA can discover if necessary and thus any deanonymization process can be properly delegated to .
Following this step, the client establishes a new hop key between themself and without revealing it to . To link this new hop key to the circuit, the client signs this new hop key with the old hop key , and sends the new key and this signature, i.e., , to . After validating the response, obtains by threshold-encrypting all now-established data.
These steps can be repeated to establish longer circuits. Only the final extension step, to the exit relay, has to be treated differently as we detail next.
@@ 5.2.3. Oblivious Authentication to the Exit Relay
In the last step, the exit relay obliviously authenticates the client. Without loss of generality, we now assume that a circuit has length three, i.e., the exit relay is . Then, receives from : the current encrypted identity , the hop key , the middle relay’s identity , and its signature . As is the last hop within the circuit, it only has to threshold-encrypt the received data one last time to yield the final encrypted identity . In case of future deanonymization, a client-circuit binding is identified by the last hop key , which we also call the client identification key, and a digest of the encrypted identity.
However, holding available encrypted identities and associated metadata of outgoing traffic flows quickly becomes expensive for the exit relay. Thus, Seldom outsources this data to government-sponsored databases. The threshold encryption of prevents unlawful access by individual parties and outsourcing this information unburdens exit relays and ensures data availability to LEA at the same time. However, exit relays only periodically upload their data in batches to mitigate traffic analysis attacks and improve bandwidth usage. The uploaded data is retained only for a predetermined period, e.g., for one year. Afterwards, the consortium and exit relays refuse to support deanonymizations.
@@ 5.3. Linking Internet Traffic to Circuits
Following an established circuit, the exit relay’s outgoing traffic must be indisputably linked to a client. Here, Seldom takes a flow-based approach. Therefore, whenever the client’s makes the exit relay establish a new flow, he signs this flow record and the database key to his encrypted identity with the client identification key. The flow record constitutes of the destination’s IP address, the targeted port, the exit relay’s IP address, and a timestamp. Only the client can generate this signature, and the corresponding public key is stored in the encrypted identity . In case of an deanonymization request, the flow record, this signature, and the encrypted identity shift the blame towards the client. It may happen that multiple similar flows by different clients are opened in short succession such that a request cannot be unambiguously linked to one encrypted identity. In these cases, LEA would have to investigate two or more circuits.
@@
In practice, when opening a new connection, the client creates the signature over the flow record and appends it to the first transmitted cell. The exit relay verifies this signature and then establishes the connection. Afterward, the client can communicate over this flow without additional overhead through the circuit. To outsource the storage of this flow record, the exit relay symmetrically encrypts the signature and the digest . The resulting data layout is displayed in Figure 3. We made the design choice to symmetrically encrypt this sensitive data and rely on the exit relay to decrypt it in case of an deanonymization request. Even if the exit relay were offline, it would be traceable through the consensus or IP address. With another threshold encryption we could avoid this involvement of the exit relay, but this would reduce oversight. Most importantly, the active involvement of exit relays ensures the correctness of the statistics and disclosure of deanonymization requests even for a dishonest consortium.
@@
An additional challenge comes from the fact that multiple governments want their own consistent copy of this database. It would put additional strain on the exit relays if they have to migrate their database entries to each government’s database. We offload these costs to the governments, while simultaneously benefiting from the better connectivity among them. However, if the governments synchronize the databases themselves, a database operator can doubt he knows about all data distributed by the exit relays. Hence, we implemented a smart contract (deployed on the translucent ledger introduced in Section 6.2) through which exit relays inform about the digest of migrated database entries. A government thus learns which data other governments should distribute to them. All in all, Seldom thus allows the storage of the necessary data for deanonymizations without overburdening the exit relays or other volunteers with computations or costs.
@@ 6. Deanonymization Process
In the following, we look at the procedure for deanonymizations, as shown in Figure 4. Seldom relies on a consortium of trusted parties to approve and execute deanonymizations. While this paper focuses on the technical aspects, we nevertheless give some considerations about the potential composition of such a consortium in Section 6.1. To ensure public oversight, we introduce translucent ledgers in Section 6.2 to manage this consortium. These ledgers allows the public to see live statistics about attempted and successful deanonymization requests, ensures the delayed information disclosure, and allows the consortium to grant selective data access to proof an approved request to an exit relay. Afterward, we discuss the deanonymization process in Section 6.3. Finally, we see how Seldom probes relays to preemptively detect rogue relays without compromising anyone’s privacy in Section 6.4.
@@ 6.1. Assembling a Trusted Consortium
We identify three roles that should be represented in Seldom’s consortium: governments, privacy advocates, and privacy-focused corporations. First, we need representatives from the judicial branch of participating governments. These entities have the knowledge to assess deanonymization requests based on the potential benefit to public safety and their legality. Secondly, the consortium needs to incorporate parties that are known to strongly advocate for people’s privacy. These parties are essential to generate trust into the system. They can block excessive requests or such requests that clearly do no benefit public safety, e.g., if the identity of a whistleblower would be requested. This group could include the operators of Tor’s directory authorities and other well-known privacy advocates. Thirdly, we see a need to include corporations that are known to prevent governments from excessively accessing encrypted data, e.g., Mozilla and Apple. These are powerful entities in comparison to typical privacy advocates and thus are better equipped to oppose political pressure.
Besides the type of representatives in the consortium, its size must also be determined, as well as the threshold to approve requests. This threshold can be lower than the threshold set for the threshold decryption, assuming that non-agreeing entities still cooperate in the decryption once a vote passed. Moreover, the interval between the renewing of threshold-encryption keys must be decided on, such that honest consortium members can ensure that no data can be access beyond its intended retention period. Overall, the consortium should be relatively static and changes planned in advance to happen with scheduled key updates, but they also require updating the smart contract. Considering the number of consortium members and the voting procedure (e.g., simple majority, majority within each role, etc.), we do not have the expertise to make a good judgement. Such questions must anyway be decided through extensive public discourse. We assume that a consortium can be assembled that at least a portion of privacy-conscious Internet users approve of.
@@ 6.2. Translucent Ledgers
To manage the consortium and guarantee public oversight, we propose translucent ledgers. In nature, these are private blockchains where new blocks (in our case containing information about deanonymization requests) are agreed upon by a consortium and cryptographically linked to the previous block. What makes these ledgers translucent is that the consortium grants access to selected data for specific clients and the public. The three pillars of translucency are: (i) delayed disclosure, (ii) peeks into real-time data, and (iii) real-time statistics. These three pillars allow to keep ledger data secret as much as necessary, while providing guarantees of the correct execution of processes to the public.
@@
Delayed Disclosure. The first building block of translucent ledgers is the notion of delayer disclosure, i.e., the delayed release of all stored information to the public such that they can be verified by everyone. For Seldom, this delayed release is based on a fixed delay, as already exists for the declassification of secret documents by democratic governments. Still, the consortium may unanimously on delaying the release of specific data, e.g., if a case is still ongoing. Here, a single entity in the consortium ensures the timely data disclosure, while the blockchain properties ensure data integrity. Importantly, this step only reveals which deanonymizations have been requested (including a reasoning) and which of those succeeded, while the client’s identity is never written to the ledger and thus also not revealed to the public.
@@
Peeks into Real-time Data. While disclosure to the public is delayed, interactions with the ledger must happen in real time. Especially exit relays must be able read the ledger when they are authorized to do so, i.e., to verify the approval of a deanonymization request. Yet, no additional information should leak about the ledger’s current state to unauthorized entities. To enable such fine-grained interactions, the operators of translucent ledgers can generate verifiable claims of the current state of parts of the ledger. We build our translucent ledger based on the Ethereum blockchain, such that we can use Merkle proofs as proposed in EIP1186 (Jentzsch and Jentzsch, 2018) to reveal the current state of each smart contract variable individually by linking it to the public header. These Merkle proofs links the revealed data to the already distributed and immutable block headers known to the public. Thus, authorized outsiders are enabled to peek into specific parts of the ledger’s state.
@@
Real-Time Statistics. Delaying information disclosure for a significant time can make a process appear opaque, making users more skeptical of its trustworthiness. To counteract this possibility, translucent ledgers can distribute real-time statistics about what is being stored and processed on the ledger. To this end, we borrow from the Bloom filters in Ethereum (Wood et al., 2019) that let clients quickly search for past events. Events are introduced by smart contract programmers and are triggered when certain conditions are met during code execution. Upon being fired, these event writes logs to the ledger. Additionally, the specific event that was triggered and its arguments are inserted into the Bloom filter of the current block header. Thus, by strategically specifying these events, a smart contract developer can control what real-time statistics are leaked into the block header.
@@
The notion of translucent ledgers thus offers a strong building block for the management of exceptional access that is neither opaque nor completely transparent to the public to not interfere with ongoing criminal investigations.

### v2

@@ 5. Oblivious Authentication
Seldomis built upon an oblivious authentication protocol integrated into onion routing. This protocol authenticates a client to an exit relay without revealing its identity. Clients have to obtain temporal IP certificates (Section 5.1) before being able to establish a circuit (Section 5.2) so that exit relays are convinced that any outgoing traffic can be irrefutably linked to the client (Section 5.3). In Appendix A, we detail how we integrate this protocol with the existing Tor handshake.
@@ 5.1. Temporal IP Certificates
Seldomneeds a way to reliably represent an individual’s identity if deanonymization is required later on. Regular, CA-issued certificates are not applicable to Seldom, as (i) obtaining them is costly and not always possible (Backes et al., 2014) and (ii) they would share excessive personal information with entry relays. Instead, we propose to identify users via temporal IP certificates, which link an ephemeral public key111Seldomuses the ed25519 signature scheme for its speed and small signatures. of a client to a timestamped IP address.222Since designing and implementing Seldom, Let’s Encrypt announced the rollout of IP Address Certificate (Gable, 2025), which would satisfy Seldom’s requirements. While not necessarily allowing recovery of a criminal’s true identity directly, deanonymizable temporal IP certificates effectively cancel out the added privacy provided by the anonymity network only for LEA bringing forward good reasons for further investigation. Only the entry relay obtains the temporal IP certificate in the clear for validation purposes. If the deanonymization reveals an invalid certificate, this serves as an indication for LEA that the client and entry relay colluded, pointing to the entry relay for further investigation.
Temporal IP certificates are issued in an automated manner by one or more CAs that are trusted by LEA. The client generates a public-private key pair for any new Internet connection (e.g., ISP-triggered reassignment of IPv4 addresses) and sends the public key to a CA for certification. In the simplest case, the CA extracts the client’s IP address from the request and creates the temporal IP certificate from that IP address, the public key, and the current timestamp. We further discuss adaptations to this process for users relying on bridge relays, e.g., for censorship evasion, in Appendix B.
@@ 5.2.1. Authentication to the Entry Relay
The purpose of our authentication protocol is to bind a temporal IP certificate to the hops within a circuit in a way that relays can only obtain information related to their direct predecessor. We hence refer to the ed25519 public keys to authenticate individual hops as hop keys .
After obtaining a valid temporal IP certificate (cf. Section 5.1), the client confidentially sends that certificate to the entry relay , which extracts the client’s first hop key from the certificate. The entry relay validates by verifying the IP certificate’s signature by the CA and ensuring that the client uses the certified IP address. Failures to validate the IP address will surface during deanonymizations; hence, is incentivized to perform this check honestly. Next, the client generates a second hop key pair (with the associated public key ) to prepare the authentication forwarding between relays and . By signing using the validated , can in turn validate the authenticity of . Finally, threshold-encrypts the client’s IP certificate, , and the signature under using a CCA-secure Threshold Public Key Encryption (TPKE) scheme. Thus, only the consortium can access the encrypted data from now on. We refer to the resulting ciphertext as the client’s encrypted identity . Our concrete implementation relies on a hybrid threshold-encryption scheme based on Shoup’s threshold RSA signatures (Shoup, 2000) and AES in GCM mode. We published the scheme as open source under the name thRSAhold and present it in more detail in Appendix C.
@@ 5.2.2. Oblivious Authentication Forwarding
Assuming that an obliviously authenticated partial circuit has been established up to relay , we now describe how to extend the authentication via another hop to . On a high level, forwards its view on the state of current authentication, and as well as its own relay identity key to . is usually an ed25519 public key and published via the relay’s descriptor (Dingledine et al., 2004). Other relays, such as unlisted bridge relays, can be identified pseudonymously via additional, Seldom-specific certificates retrieved from a bridge authority. Only in cases where LEA must identify a relay operator, bridge authorities would reveal the certificate owner. further signs the transmitted information with its identity key, yielding the signature . This signature is transferred to, and verified by, . Verifying the identity of reassures that LEA can discover if necessary, and thus any deanonymization process can be properly delegated to .
Following this step, the client establishes a new hop key between itself and without revealing it to . To link this new hop key to the circuit, the client signs this new hop key with the old hop key and sends the new key and this signature, i.e., , to . This signature is the only link between two hop keys. To prevent collusion attacks between relays, it is, therefore, crucial that this signature is only revealed to the relay also knowing the corresponding keys, i.e., . After validating the signature, computes by threshold-encrypting all now-established data.
These steps can be repeated to establish longer circuits. Only the final extension step, to the exit relay, has to be treated differently, as we detail next.
@@ 5.2.3. Oblivious Authentication to the Exit Relay
In the last step, the exit relay obliviously authenticates the client. Without loss of generality, we now assume that a circuit has length three, i.e., the exit relay is . Then, receives the following from : the current encrypted identity , the hop key , the middle relay’s identity , and its signature . As is the last hop within the circuit, it only has to threshold-encrypt the received data one last time to yield the final encrypted identity . In case of future deanonymization, a client-circuit binding is identified by the last hop key , which we also call the client identification key, and a digest of the encrypted identity (Seldom uses SHA-3 as its cryptographic hash algorithm ).
However, storing available encrypted identities and associated metadata of outgoing traffic flows quickly becomes expensive for the exit relay. Thus, Seldom outsources this data to external databases. The threshold encryption of prevents unlawful access by individual parties and outsourcing this information unburdens exit relays and ensures data availability to LEA at the same time. Therefore, exit relays are expected to upload their data periodically in batches to mitigate traffic analysis attacks and improve bandwidth usage. The uploaded data is retained only for a predetermined period, e.g., for one year. Afterward, the consortium and exit relays refuse to support deanonymizations.
@@ 5.3. Linking Internet Traffic to Circuits
Following an established circuit, the exit relay’s outgoing traffic must be indisputably linked to a client. Here, Seldom takes a flow-based approach. Therefore, whenever the client makes the exit relay establish a new flow, it signs this flow record and the database key to its encrypted identity with the client identification key. The flow record, illustrated in Figure 3, consists of the destination’s IP address, the targeted port, the exit relay’s IP address, and a timestamp. Only the client can generate this signature, and the corresponding public key is stored in the encrypted identity . In case of a deanonymization request, the flow record, this signature, and the encrypted identity shift the blame toward the client.
@@
In practice, when opening a new connection, the client creates the signature over the flow record and includes it in the first transmitted cell. The exit relay verifies this signature and then establishes the connection. Afterward, the client can communicate over this flow without additional overhead through the circuit. To outsource the storage of this flow record, the exit relay symmetrically encrypts the signature and the digest . The resulting data layout is displayed in Figure 3. We made the design choice to symmetrically encrypt this sensitive data and rely on the exit relay to decrypt it in case of a deanonymization request. Even if the exit relay were offline, it would be traceable through the consensus or IP address. With another threshold encryption, we could avoid this involvement of the exit relay, but this would reduce oversight. Most importantly, the active involvement of exit relays ensures the correctness of the statistics and disclosure of deanonymization requests even for a dishonest consortium.
In theory, two different clients could open a connection through the same exit relay in close succession. In this case, it would be hard to match observed Internet traffic to a specific flow record. This challenge stems from the fact that the client cannot reliably know the port used by an exit relay; hence, the client cannot include this information in its signature. In practice, we expect such similar flow records to rarely and mostly happen to widely known and benign websites, e.g., news sources. In the worst case, LEA would have to investigate two or more circuits, which could come with higher scrutiny by the consortium deciding on the deanonymization request.
@@
An additional challenge comes from the fact that no single database operator is trusted by everybody. It would, however, put additional strain on the exit relays if they have to migrate their database entries to multiple deployed databases. We offload these synchronization costs to the operators while simultaneously benefiting from the better connectivity among them. However, if the operators synchronize the databases themselves, one operator must be sure that they know about all data distributed by the exit relays. Hence, we implemented a smart contract (deployed on the translucent ledger introduced in Section 6.2) through which exit relays inform about the digest of migrated database entries. Each database operator thus learns which data other instances should distribute to them. All in all, Seldom thus allows the distributed storage of the necessary data for deanonymizations without overburdening the exit relays or other volunteers with additional computations or costs.
5.4. A Seldom Circuit from the Perspective of an Exit Relay
[…]
@@ 6. Deanonymization Process
In the following, we look at Seldom’s deanonymization process, which is illustrated by Figure 5. Seldom relies on a consortium of trusted parties to approve and execute deanonymizations. While this paper focuses on the technical aspects, we nevertheless give some considerations about the potential composition of such a consortium in Section 6.1. To ensure public oversight, we introduce translucent ledgers in Section 6.2 to manage this consortium. The use of a translucent ledger enables the public to see live statistics about attempted and successful deanonymization requests, ensures the delayed information disclosure, and allows the consortium to grant selective data access to prove an approved request to an exit relay. Afterward, we discuss the deanonymization process in Section 6.3. Finally, we see how Seldom probes relays to preemptively detect rogue relays without compromising anyone’s privacy in Section 6.4.
@@ 6.1. Assembling a Trusted Consortium
We identify three roles that should be represented in Seldom’s consortium: governments, privacy advocates, and privacy-focused corporations. First, we need representatives from the judicial branch of participating governments. These entities have the knowledge to assess deanonymization requests based on the potential benefit to public safety and their legality. Secondly, the consortium needs to incorporate parties that are known to strongly advocate for people’s privacy. These parties are essential to generate trust in the system. They can block excessive requests or such requests that clearly do not benefit public safety, e.g., if the deanonymization of a whistleblower would be requested. This group could include the operators of Tor’s directory authorities and other well-known privacy advocates. Thirdly, we see a need to include corporations that are known to prevent governments from excessively accessing encrypted data, e.g., Mozilla and Apple. These are powerful entities relative to typical privacy advocates and thus are better equipped to oppose political pressure.
Besides the type of representatives in the consortium, its size must also be determined, as well as the threshold to approve requests. This threshold can be lower than the threshold set for the threshold decryption, assuming that non-agreeing entities still cooperate in the decryption once a vote has passed. Moreover, the interval between the renewing of threshold-encryption keys must be decided on, such that honest consortium members can ensure that no data can be accessed beyond its intended retention period. Overall, the consortium should be relatively static, and changes should be planned in advance to happen with scheduled key updates, but they also require updating the smart contract. Considerations regarding the consortium’s size and the voting procedure (e.g., simple majority, majority within each role, etc.) remain important aspects to Seldom’s deployment but go beyond the scope of this work.
@@ 6.2. Translucent Ledgers
We propose the concept of translucent ledgers for the coordination among the consortium while ensuring public oversight. Generally, a translucent ledger is a private blockchain where a new block (deanonymization requests in Seldom’s case) is accepted by the consensus of the consortium members and is cryptographically linked to its predecessor. For Seldom, we use an Ethereum (Wood and others, 2019) blockchain with the Proof of Authority (PoA) consensus mechanism, which requires minimal modifications to be used as the translucent ledger. What makes a ledger translucent is that the consortium grants access to selected data for specific clients and the public. The three pillars of translucency are (i) delayed disclosure, (ii) peeks into real-time data, and (iii) real-time statistics. These three pillars allow keeping ledger data secret as much as necessary while providing guarantees of the correct execution of processes to the public.
@@
Delayed Disclosure. The first building block of translucent ledgers is the notion of delayed disclosure, i.e., the delayed release of all stored information to the public such that they can be verified by everyone. For Seldom, this delayed release is based on a fixed delay, as already exists for the declassification of secret documents by democratic governments. Still, the consortium may unanimously agree on delaying the release of specific data, e.g., if a case is still ongoing. Here, a single honest entity in the consortium suffices to enforce the timely data disclosure, while the blockchain properties ensure data integrity. Importantly, this step only reveals which deanonymizations have been requested (including a reasoning) and which of those succeeded, while the client’s identity is never written to the ledger and thus also not revealed to the public.
@@
Peeks into Real-time Data. While disclosure to the public is delayed, interactions with the ledger must happen in real time. Especially exit relays must be able to read the ledger when they are authorized to do so, i.e., to verify the approval of a deanonymization request. Yet, no additional information should leak about the ledger’s current state to unauthorized entities. To enable such fine-grained interactions, the operators of translucent ledgers can generate verifiable claims of the current state of parts of the ledger. We build our translucent ledger based on the Ethereum blockchain, such that we can use Merkle proofs as proposed in EIP1186 (Jentzsch and Jentzsch, 2018) to reveal the current state of each smart contract variable individually by linking it to the public header. These Merkle proofs link the revealed data to the already distributed and immutable block headers known to the public. Thus, authorized outsiders are enabled to peek into specific parts of the ledger’s state.
@@
Real-Time Statistics. Delaying information disclosure for a significant time can make a process appear opaque, making users more skeptical of its trustworthiness. To counteract this possibility, translucent ledgers can distribute real-time statistics about what is being stored and processed on the ledger. To this end, we use the Bloom filters in Ethereum (Wood and others, 2019) that usually are used to let clients quickly search for past events. Events are introduced by smart contract programmers and are triggered when certain conditions are met during code execution. Upon being fired, these events write logs to the ledger. Additionally, the specific event that was triggered and its arguments are inserted into the Bloom filter of the current block header. Thus, by strategically specifying these events, a smart contract developer can control what real-time statistics are leaked into the block header.
@@
Translucent ledgers are thus a strong building block for the management of exceptional access that is neither opaque nor completely transparent to the public.